OpenStudy (anonymous):

The mean of four consecutive even numbers is 15. The greatest of these numbers is ? The least of these numbers is ?

OpenStudy (anonymous):

x+ (x+2) + (x+4) + (x+6) =15

OpenStudy (anonymous):

find the value of x. x+6 is the greatest x is lowest

OpenStudy (anonymous):

so 6 is the greatest and x is the lowest?

OpenStudy (anonymous):

x+6 is the greatest...first u have to find the value of x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and how do i do that

OpenStudy (amistre64):

..by adding it all up to start with

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

after that, subtract and divide ....

OpenStudy (anonymous):

x+ (x+2) + (x+4) + (x+6) =15 <-------------------- so add all that up

OpenStudy (amistre64):

the left side of the equals sgn ... yes

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

and i see a slight typing error; the mean of 4 numbers is their sum divided by 4

OpenStudy (amistre64):

\[\frac{x+x+2+x+4+x+6}{4}=15\]

OpenStudy (anonymous):

i dont get it

OpenStudy (amistre64):

i dont read minds sooo ....

geerky42 (geerky42):

Do you know what mean is?